Summary of LASCO/EIT Recommissioning Steps

Recommissioning began Tuesday October 13 at 1:00 PM EST (17:00 UT)

EIT images have been successfully taken in all four wavelengths.

C3 Coronal images have been taken and appear fine.

C2 Coronal images have been taken and appear fine.

Update - Wednesday October 15th: The C2 door has been opened and C2 coronal images have been taken. See latest images here (note: the C3 images have not had the f-corona/stray-light model subtracted because the spacecraft is rolled ~53 degrees. It will probably be rotated back on Friday to put solar north up.)

A more detailed status report is availaible as a separate file from the normal status reports.

This is a summary of the steps that will be followed during the recommissioning of the LASCO and EIT instruments. A check mark will appear when the steps are completed along with the date and time of completion. After Block 2, the blocks could be performed in a different order than what appears here.

Block 1
Boot all processors, turn on thermal control, check proper interprocess communications

Block 2
Check EIT mechanisms and camera, Image XUV Corona, Check C2/C3 mechanisms, camera

Block 3
Open C3 Door, First Light Image of C3 Corona

Block 4
Open C2 Door, First Light Image of C2 Corona

Block 5
Door Closed Calibration of C3

Block 6
Door Closed Calibration of C2



Block 7
EIT Calibration

Block 8
Check out C1 mechanisms, camera and Fabry-Perot

Block 9
Re-open C2/C3 Doors, calibration checks of C2/C3